OpenStudy (anonymous):

can someone explain this one to me? If f(2)= 10, f(3) = 20, f(4) = 30, f(5) = 40, then f(7) =

OpenStudy (anonymous):

f ( 7) - f(2) / ( 7 -2 ) = 10 -> f (7) - f(2) = 50 =>f(7) = 50 + 10 = 60